# Canonical Representation and Force-Based Pretraining of 3D Tactile for Dexterous Visuo-Tactile Policy Learning

### v2

@@ Abstract
Tactile sensing plays a vital role in enabling robots to perform fine-grained, contact-rich tasks. However, the high dimensionality of tactile data, due to the large coverage on dexterous hands, poses significant challenges for effective tactile feature learning, especially for 3D tactile data, as there are no large standardized datasets and no strong pretrained backbones. To address these challenges, we propose a novel canonical representation that reduces the difficulty of 3D tactile feature learning and further introduces a force-based self-supervised pretraining task to capture both local and net force features, which are crucial for dexterous manipulation. Our method achieves an average success rate of 78% across four fine-grained, contact-rich dexterous manipulation tasks in real-world experiments, demonstrating effectiveness and robustness compared to other methods. Further analysis shows that our method fully utilizes both spatial and force information from 3D tactile data to accomplish the tasks. The videos can be viewed at https://3dtacdex.github.io.
@@ I INTRODUCTION
Considering the power of visual backbones, many works [13, 4] convert tactile data into 2D images to reduce the complexity of learning useful tactile features. However, this transformation leads to the change and loss of part spatial information between different taxels. To preserve these spatial relationships, most approaches represent 3D tactile data as a graph and use graph neural networks (GNNs)[14] to encode tactile signals[12, 15]. However, these methods primarily focus on specific tasks and require a large amount of data to learn effective features. Inspired by the success of the pretraining strategy in vision-based learning, T-DEX [4] collects tactile play data through interaction with various objects and pretrain tactile encoder with self-supervised learning. This pretraining improves the efficiency of feature learning and enhances diverse downstream robotic manipulation tasks. However, it still relies on 2D images as the tactile representation. As a result, efficiently learning 3D tactile data features for dexterous manipulation remains a challenge.
Figure 1: Real Robot System. Our system uses a single camera and distributed tactile sensors to achieve dexterous, fine-grained, contact-rich tasks. Note that the teleoperation camera is only used for data collection, not policy learning.
@@ IV METHOD
We focus on the problem of leveraging 3D tactile data from distributed tactile sensors for learning visuo-tactile dexterous manipulation policies. To reduce the difficulty of learning features from complex 3D tactile data, we canonicalize the data into a unit frame, as described in IV-A. We then pretrain the tactile encoder using self-supervised force-based prediction tasks to enhance local and net force feature learning, detailed in IV-B. This pretrained tactile encoder is subsequently used for visuo-tactile policy learning, as outlined in IV-C.
@@ IV-B Force-based Pretraining
What kind of pretraining can we use for 3D tactile data to improve dexterous manipulation policies? When humans manipulate objects, we carefully apply force to achieve the desired object pose. Inspired by this, we propose pretraining the 3D tactile data based on force. When applying force to an object, it is essential to consider how each part of the finger applies local force so that the net force moves the object as intended. Consequently, we designed two force-based self-supervised pretraining tasks: the first predicts the local force, and the second predicts the net force. Since our robot system differs from T-DEX [4], we follow their method to collect our own play data for pretraining. We use GraphMAE [32] as the backbone for pretraining.
@@
To help the encoder learn the local features of each taxel, we design a masked force prediction task. Since the force applied to each taxel can propagate to its neighboring taxels, we randomly mask part of the tactile force in the play data and use the masked tactile data as input for the encoder. The encoder first encodes the tactile data, then decodes the latent representation to reconstruct the original tactile data, as shown in Fig. 2. We use the MSE loss to compute the loss between the reconstructed and original force values, only for the masked forces. This pretraining approach helps the encoder learn the relationships between local forces.
@@ V-B Baselines
We compare our method with the following baselines, which all use the same visual backbone, diffusion policy backbone, visual observation, robot proprioceptive state, and action space as ours, but with different types of tactile representation and pertaining. 1) DP: We implement the diffusion policy without using tactile data or pertaining for this baseline. 2) HATO: HATO [20] uses MLP to directly encode the tactile readings. We flatten 360-dimensional force values and use MLP to encode tactile data for this baseline. 3) T-DEX: T-DEX [4] convert the raw tactile readings into 2D image, and pretrain with BYOL [27]. For this baseline, we follow their procedure, first converting raw, tactile readings into 2D images, then pretrain on our own collected dataset and we use the pretrained encoder for diffusion policy learning. 4) GNN: We use the 9D tactile representation (e.g., the 6D pose of each taxel with respect to the hand’s base with 3D tactile force values) as input for this baseline, use graph attention networks [38] to encode tactile, which is the same GNN backbone as ours. Since there are no pertaining strategy designed especially for 9D tactile representation, We do not pretrain for this baseline.
@@ V-D Importance of Representation and Pretraining
To validate the effectiveness of our canonical representation and force-based pretraining, we conduct ablation studies across all tasks. As shown in Table II, using the canonical representation achieves a 48% success rate, even without pretraining. However, without the canonical representation, even with pretraining, the policy fails to complete any task. We observed that the policy directly moves to a specific hand joint position upon starting inference and then repeats similar actions. Analyzing the output of tactile encoder, we found that without the canonical representation, the encoder outputs similar features for taxels within the same sensor pad, failing to perceive fine-grained differences. This validates the necessity of the canonical representation. Based on such representation, pretraining further enables the encoder to learn more useful features, increasing the success rate to 78%.
@@ VI CONCLUSIONS
In this work, we enhance 3D tactile feature learning by proposing a novel canonical representation that aligns differently distributed tactile sensor readings, reduces the feature space, and increases the discriminability of each taxel within the same sensor. We also introduce a force-based self-supervised pretraining task to encourage the use of both spatial and force information. Real-world experiments using the pretrained encoder for downstream dexterous, fine-grained, contact-rich tasks demonstrate the effectiveness and robustness of our methods.

### v3

@@ Abstract
Tactile sensing plays a vital role in enabling robots to perform fine-grained, contact-rich tasks. However, the high dimensionality of tactile data, due to the large coverage on dexterous hands, poses significant challenges for effective tactile feature learning, especially for 3D tactile data, as there are no large standardized datasets and no strong pretrained backbones. To address these challenges, we propose a novel canonical representation that reduces the difficulty of 3D tactile feature learning and further introduces a force-based self-supervised pretraining task to capture both local and net force features, which are crucial for dexterous manipulation. Our method achieves an average success rate of 78% across four fine-grained, contact-rich dexterous manipulation tasks in real-world experiments, demonstrating effectiveness and robustness compared to other methods. Further analysis shows that our method fully utilizes both spatial and force information from 3D tactile data to accomplish the tasks. The codes and videos can be viewed at https://3dtacdex.github.io.
@@ I INTRODUCTION
Considering the power of visual backbones, many works [13, 4] convert tactile data into 2D images to reduce the complexity of learning useful tactile features. However, this transformation leads to the change and loss of part spatial information between different taxels. To preserve these spatial relationships, most approaches represent 3D tactile data as a graph and use graph neural networks (GNNs)[14] to encode tactile signals[12, 15]. However, these methods focus on specific tasks and require large data to learn effective features. Inspired by the success of the pretraining strategy in vision-based learning, T-DEX [4] collects tactile play data through interaction with various objects and pretrain tactile encoder with self-supervised learning. This pretraining improves the efficiency of feature learning and enhances diverse downstream robotic manipulation tasks. However, it still relies on 2D images as the tactile representation. As a result, efficiently learning 3D tactile data features for dexterous manipulation remains a challenge.
Figure 1: Real Robot System. Our system uses one camera and distributed tactile sensors to achieve dexterous, fine-grained, contact-rich tasks. The teleoperation camera is only used for data collection, not policy learning.
@@ IV METHOD
We focus on the problem of leveraging 3D tactile data from distributed tactile sensors for learning visuo-tactile dexterous manipulation policies. To reduce the difficulty of learning features from complex 3D tactile data, we canonicalize the data into a unit frame, in IV-A. We then pretrain the tactile encoder using self-supervised force-based prediction tasks to enhance local and net force feature learning, in IV-B. This pretrained tactile encoder is subsequently used for visuo-tactile policy learning, in IV-C.
@@ IV-B Force-based Pretraining
What kind of pretraining can we use for 3D tactile data to improve dexterous manipulation policies? When humans manipulate objects, we carefully apply force to achieve the desired object pose. Inspired by this, we propose pretraining the 3D tactile data based on force. When applying force, it is essential to consider how each finger part applies local force so that the net force moves the object as intended. Consequently, we designed two force-based self-supervised pretraining tasks: the first predicts the local force, and the second predicts the net force. Since our robot system differs from T-DEX [4], we follow their method to collect our own play data for pretraining. We use GraphMAE [32] as the backbone for pretraining.
@@
To help the encoder learn the local features of each taxel, we design a masked force prediction task. Since the force applied to each taxel can propagate to its neighboring, we randomly mask part of the tactile force and use the masked tactile data as input for the encoder. The encoder first encodes the tactile data, then decodes the latent representation to reconstruct the original tactile data, as shown in Fig. 2. We compute MSE loss between the reconstructed and original force values, only for the masked forces. This pretraining approach helps the encoder learn the relationships between local forces.
@@ V-B Baselines
We compare our method with the following baselines, which all use the same visual backbone, diffusion policy backbone, visual observation, robot proprioceptive state, and action space as ours, but with different types of tactile representation and pertaining. 1) DP: We implement the diffusion policy without using tactile data or pertaining for this baseline. 2) HATO: HATO [20] uses MLP to encode the tactile. We flatten force values and use MLP to encode tactile data for this baseline. 3) T-DEX: T-DEX [4] convert the raw tactile into 2D image, and pretrain with BYOL [27]. For this baseline, we follow their procedure, first converting raw tactile into 2D images, then pretrain encoder on our own collected dataset, then encoder for diffusion policy learning. 4) GNN: We use the 9D tactile representation (e.g., the 6D pose of each taxel with 3D tactile force values) as input for this baseline, use graph attention networks [38] to encode tactile, which is the same GNN backbone as ours. Since there are no pertaining strategy designed especially for 9D tactile representation, We do not pretrain for this baseline.
@@ V-D Importance of Representation and Pretraining
To validate the effectiveness of our canonical representation and force-based pretraining, we conduct ablation studies across all tasks. As shown in Tab. II, using canonical representation achieves 48% success rate, even without pretraining. However, without canonical representation, even with pretraining, the policy fails to complete any task. We observed that the policy moves to a specific hand joint position upon starting inference and then repeats similar actions. Analyzing the output of tactile encoder, we found that without canonical representation, the encoder outputs similar features for taxels within the same sensor pad, failing to perceive fine-grained differences. This validates the necessity of canonical representation. Based on such representation, pretraining further enables the encoder to learn more useful features, increasing the success rate to 78%.
@@ VI CONCLUSIONS
In this work, we enhance 3D tactile feature learning by proposing a novel canonical representation that aligns differently distributed tactile sensor readings, reduces the feature space, and increases the discriminability of each taxel within the same sensor. We also introduce a force-based self-supervised pretraining task to encourage using both spatial and force information. Real-world experiments using the pretrained encoder for downstream dexterous, fine-grained, contact-rich tasks demonstrate the effectiveness and robustness of our methods.